# A New Algorithm for Zero-Sum Linear-Quadratic Stochastic Differential Games in Infinite Horizons

### v1

@@ 3.1 Analysis of Structural Characteristics for ZSLQSDG
  • $0\in\bm{\Gamma}^{[A,C_{1},\dots,C_{r};B,D_{1},\dots,D_{r};Q,R,S]}$.
  • The system $A,C_{1},\dots,C_{r};\ B,D_{1},\dots,D_{r}]$ is mean-square stabilizable.
  • There exists a matrix set $\{E_{0},E_{1},\dots,E_{r}\}$ satisfying $\sum_{l=0}^{r}E_{l}^{\top}E_{l}=Q$ such that the system
    $E_{0},E_{1},\dots,E_{r};A,C_{1},\dots,C_{r}]$
@@
By Theorem 4.7 in [13], along with the condition $0\in\bm{\Gamma}$ and the mean-square stabilizability of $\left[A,C_{1},\dots,C_{r};B,D_{1},\dots,D_{r}\right],$ the ARE (16) has a maximal solution $P_{\text{max}}\in\overline{\mathbb{S}}_{+}^{n}$ such that $R+\sum_{l=1}^{r}D_{l}^{\top}P_{\text{max}}D_{l}\succ 0$. Using Proposition 3.2, the ARE for $P_{\text{max}}$ can be rewritten as:
@@
Since the system $\left[E_{0},E_{1},\dots,E_{r};A,C_{1},\dots,C_{r}\right]$ is stochastically detectable, there exists $\varTheta\in\mathbb{R}^{n\times q}$ such that $\left(A+\varTheta E_{0},C_{1}+\varTheta E_{1},\dots,C_{r}+\varTheta E_{r}\right)$ is mean-square stable. Define
$\hat{\varTheta}=\begin{bmatrix}\varTheta^{\times(r+1)}&-B\sqrt{(r+1)R^{-1}}&-D_{1}\sqrt{(r+1)R^{-1}}&\cdots&-D_{r}\sqrt{(r+1)R^{-1}}\end{bmatrix}\in\mathbb{R}^{n\times[(q+m)(r+1)]},$
then the system $\left(A+BT(P_{\text{max}})+\hat{\varTheta}\hat{E}_{0},C_{1}+D_{1}T(P_{\text{max}})+\hat{\varTheta}\hat{E}_{0},\dots,C_{r}+D_{r}T(P_{\text{max}})+\hat{\varTheta}\hat{E}_{0}\right)$ is mean-square stable. By Lemma 3.5, the system $(A+BT(P_{\text{max}}),\ C_{1}+D_{1}T(P_{\text{max}}),\ \dots,\ C_{r}+D_{r}T(P_{\text{max}}))$ is also mean-square stable. Hence, $P_{\text{max}}$ is a stabilizing solution of ARE (16). By Theorem 5.6.5 in [12], the stabilizing solution is unique, so ARE (16) has a unique stabilizing solution $P\in\overline{\mathbb{S}}_{+}^{n}$ such that $R+\sum_{l=1}^{r}D_{l}^{\top}PD_{l}\succ 0$. ∎
@@ 3.3 Convergence Analysis of Iterative Sequence
Next, we introduce two linear operators associated with the interrelated iterative steps, whose specific properties are analyzed in Lemma 3.8. These favorable properties are crucial for proving the convergence of the algorithm’s iterative process, as they jointly ensure the boundedness of the iterative sequence—a fact explicitly demonstrated in the proof of Theorem 3.9.
@@
$\mathcal{L}^{*}_{P}(Y)=$ $Y\left(A_{(0)}+B_{1}\hat{K}_{1}(P)+B_{2}L\right)+\left(A_{(0)}+B_{1}\hat{K}_{1}(P)+B_{2}L\right)^{\top}Y$ (19)
$+\sum_{l=1}^{r}\left(C_{l,(0)}+D_{l,1}\hat{K}_{1}(P)+D_{l,2}L\right)^{\top}Y\left(C_{l,(0)}+D_{l,1}\hat{K}_{1}(P)+D_{l,2}L\right)\,,\forall\,Y\in\mathbb{S}^{n};$
$\mathcal{L}^{*}_{P+Z}(Y)=$ $Y\left(A_{(0)}+B_{1}\hat{K}_{1}(P+Z)+B_{2}L\right)+\left(A_{(0)}+B_{1}\hat{K}_{1}(P+Z)+B_{2}L\right)^{\top}Y$ (20)
$+\sum_{l=1}^{r}\left(C_{l,(0)}+D_{l,1}\hat{K}_{1}(P+Z)+D_{l,2}L\right)^{\top}Y\left(C_{l,(0)}+D_{l,1}\hat{K}_{1}(P+Z)+D_{l,2}L\right)\,,\forall\,Y\in\mathbb{S}^{n},$
where
@@
Assume $0\in\mathrm{Dom}\,\mathcal{G}$, there exists an $L\in\mathcal{A}$ and $P,Z\in\mathbb{S}^{n}$ having the properties:
@@
    $\quad\times[B_{2}^{\top}Z+\sum_{l=1}^{r}D_{l,2}^{\top}Z(C_{l,(0)}+D_{l,1}\hat{K}_{1}(P)+D_{l,2}\hat{K}_{2}(P))]=0.$
Let $L\in\mathcal{A}$ and $\tilde{P}_{L}$ be the stabilizing solution of the ARE (7). If $\mathcal{L}^{*}_{P}$ , $\mathcal{L}^{*}_{P+Z}$ are linear operators associated to the triple $(P,Z,L)$ via (19) and (20), then the following assertions hold:
If   $\text{Spec}\,\mathcal{L}^{*}_{P}\subset\mathbb{C}^{-}$, then $\tilde{P}_{L}\succeq P+Z$.
If   $\tilde{P}_{L}\succeq P+Z$, then $\text{Spec}\,\mathcal{L}^{*}_{P+Z}\subset\mathbb{C}^{-}$.
@@
First, we prove the Conclusion (i.). From Proposition 3.2, we can obtain:
@@
where $J(P)=\begin{bmatrix}\hat{K}_{1}(P)\\ L\end{bmatrix}$, $\mathcal{L}^{*}_{P}$ is defined in (19), $\hat{K}_{1}(P)$ and $\hat{K}(P)$ is defined in (21). Combining Eq. (22) and Proposition 3.3, we find that $P+Z$ solves the following ARE:
@@
where $\hat{N}_{1}(P,Z)=B_{1}^{\top}Z+\sum_{l=1}^{r}D_{l,1}^{\top}Z(C_{l,(0)}+D_{l,1}\hat{K}_{1}(P))+D_{l,2}\hat{K}_{2}(P))$, $\hat{N}_{2}(P,Z)=B_{2}^{\top}Z+\sum_{l=1}^{r}D_{l,2}^{\top}Z(C_{l,(0)}+D_{l,1}\hat{K}_{1}(P))+D_{l,2}\hat{K}_{2}(P))$ and $\mathbb{R}_{22}^{\sharp}(P+Z)$ is defined in (12).
Moreover,
$(J(P)-\hat{K}(P+Z))^{\top}R(P+Z)(J(P)-\hat{K}(P+Z))$ (24)
$=(\hat{K}_{1}(P)-\hat{K}_{1}(P+Z))^{\top}\mathbb{R}_{22}^{\sharp}(P+Z)(\hat{K}_{1}(P)-\hat{K}_{1}(P+Z))+H^{\top}_{1}H_{1}=$
$(\hat{N}_{1}(P,Z)-R_{12}(P+Z)R_{22}(P+Z)^{-1}\hat{N}_{2}(P,Z))^{\top}\mathbb{R}_{22}^{\sharp}(P+Z)^{-1}(\hat{N}_{1}(P,Z)-R_{12}(P+Z)R_{22}(P+Z)^{-1}\hat{N}_{2}(P,Z))$
$+H_{1}^{\top}H_{1},$
where $H_{1}=\begin{bmatrix}R_{22}(P+Z)^{-\frac{1}{2}}R_{21}(P+Z)&R_{22}(P+Z)^{\frac{1}{2}}\end{bmatrix}(J(P)-\hat{K}(P+Z))$. Next, by virtue of Proposition 3.2, $\tilde{P}_{L}$ can be rewritten as:
@@
$-(\hat{K}_{1}(P)-K_{L}(\tilde{P}_{L}))^{\top}(R_{11}+\sum_{l=1}^{r}D_{l,1}^{\top}\tilde{P}_{L}D_{l,1})(\hat{K}_{1}(P)-K_{L}(\tilde{P}_{L}))=0.$
where $K_{L}(\tilde{P}_{L})=-(R_{11}+\sum_{l=1}^{r}D_{l,1}^{\top}\tilde{P}_{L}D_{l,1})^{-1}(B^{\top}_{L}\tilde{P}_{L}+\sum_{l=1}^{r}D^{\top}_{l,1}\tilde{P}_{L}C_{l}+R_{12}L)$.
Subtracting Eq. (23) from Eq. (25) and combining it with Eq. (24) yields:
@@
Since the $\text{Spec}\,\mathcal{L}^{*}_{P}\subset\mathbb{C}^{-}$ and $H^{\top}H\succeq 0$, by the similarly proof of (a.) $\Rightarrow$ (b.) in Lemma 3.5, we have $\tilde{P}_{L}\succeq P+Z$. This completes the proof of Conclusion (i.).
Next, we proceed to prove Conclusion (ii.). Similarly, by Proposition 3.2 and 3.3, and through combining Eq. (13) with Eq. (22), we can derive that $P+Z$ is a solution to the following ARE:
$\mathcal{L}^{*}_{P+Z}(P+Z)+Q_{L}+\hat{K}^{\top}_{1}(P+Z)R_{12}L+L^{\top}R_{21}\hat{K}_{1}(P+Z)+\hat{K}_{1}^{\top}(P+Z)R_{11}\hat{K}_{1}(P+Z)$- (26)
$(J(P+Z)-\hat{K}(P+Z))^{\top}R(P+Z)(J(P+Z)-\hat{K}(P+Z))+(\hat{N}_{1}(P,Z)-R_{12}(P+Z)R_{22}(P+Z)^{-1}\hat{N}_{2}(P,Z))^{\top}$
$\mathbb{R}_{22}^{\sharp}(P+Z)^{-1}(\hat{N}_{1}(P,Z)-R_{12}(P+Z)R_{22}(P+Z)^{-1}\hat{N}_{2}(P,Z))=0.$
@@
$\mathcal{L}^{*}_{P+Z}(\tilde{P}_{L})+Q_{L}+\hat{K}^{\top}_{1}(P+Z)R_{12}L+L^{\top}R_{21}\hat{K}_{1}(P+Z)+\hat{K}_{1}^{\top}(P+Z)R_{11}\hat{K}_{1}(P+Z)$ (27)
$-(\hat{K}_{1}(P+Z)-K_{L}(\tilde{P}_{L}))^{\top}(R_{11}+\sum_{l=1}^{r}D_{l,1}^{\top}\tilde{P}_{L}D_{l,1})(\hat{K}_{1}(P+Z)-K_{L}(\tilde{P}_{L}))=0.$
@@
$(\hat{K}_{1}(P+Z)-\hat{K}_{1}(P+Z))^{\top}\mathbb{R}_{22}^{\sharp}(P+Z)(\hat{K}_{1}(P+Z)-\hat{K}_{1}(P+Z))+H^{\top}_{2}H_{2}=H^{\top}_{2}H_{2}$ (28)
where $H_{2}=\begin{bmatrix}R_{22}(P+Z)^{-\frac{1}{2}}R_{21}(P+Z)&R_{22}(P+Z)^{\frac{1}{2}}\end{bmatrix}(J(P+Z)-\hat{K}(P+Z))$. Subtracting Eq. (26) from Eq. (28) combining it with Eq. (27) yields:
$\mathcal{L}^{*}_{P+Z}(\tilde{P}_{L}-P-Z)-(\hat{K}_{1}(P+Z)-K_{L}(\tilde{P}_{L}))^{\top}(R_{11}+\sum_{l=1}^{r}D_{l,1}^{\top}\tilde{P}_{L}D_{l,1})(\hat{K}_{1}(P+Z)-K_{L}(\tilde{P}_{L}))$
$+H_{2}^{\top}H_{2}-(\hat{N}_{1}(P,Z)-R_{12}(P+Z)R_{22}(P+Z)^{-1}\hat{N}_{2}(P,Z))^{\top}\mathbb{R}_{22}^{\sharp}(P+Z)^{-1}$
@@
Let $\Delta=\tilde{P}_{L}-P-Z$, we have: $\mathcal{L}^{*}_{P+Z}\Delta+\sum_{l=0}^{r}\hat{E}_{l}^{\top}\hat{E}_{l}=0,\Delta\succeq 0$, where ($0\leq l\leq r$)
@@
$\hat{E}_{l}=\begin{bmatrix}\mathbb{O}_{m_{1}\times n}^{\times(l)}\\ \sqrt{-\frac{1}{r+1}(R_{11}+\sum_{l=1}^{r}D_{l,1}^{\top}\tilde{P}_{L}D_{l,1})}(\hat{K}_{1}(P+Z)-K_{L}(\tilde{P}_{L}))\\ \mathbb{O}_{m_{1}\times n}^{\times(r-l)}\\ \frac{1}{\sqrt{r+1}}H_{2}\\ \sqrt{-\frac{1}{r+1}\mathbb{R}_{22}^{\sharp}(P+Z)^{-1}}(\hat{N}_{1}(P,Z)-R_{12}(P+Z)R_{22}(P+Z)^{-1}\hat{N}_{2}(P,Z))\end{bmatrix}\in\mathbb{R}^{[m_{2}+m_{1}(r+2)]\times n}.$
Choose
@@
$\hat{\varTheta}=\begin{bmatrix}-\sqrt{-(r+1)(R_{11}+\sum_{l=1}^{r}D_{l,1}^{\top}\tilde{P}_{L}D_{l,1})^{-1}}^{\top}B_{1}^{\top}\\ -\sqrt{-(r+1)(R_{11}+\sum_{l=1}^{r}D_{l,1}^{\top}\tilde{P}_{L}D_{l,1})^{-1}}^{\top}D_{1,1}^{\top}\\ \dotsb\\ -\sqrt{-(r+1)(R_{11}+\sum_{l=1}^{r}D_{l,1}^{\top}\tilde{P}_{L}D_{l})^{-1}}^{\top}D_{r,1}^{\top}\\ \mathbb{O}_{n\times m_{2}}^{\top}\\ \mathbb{O}_{n\times m_{1}}^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{n\times[m_{2}+m_{1}(r+2)]}.$
@@
We obtain that $(A_{(0)}+B_{1}\hat{K}_{1}(P+Z)+B_{2}L+\hat{\varTheta}\hat{E}_{0},C_{1,(0)}+D_{1,1}\hat{K}_{1}(P+Z)+D_{1,2}L+\hat{\varTheta}\hat{E}_{1},\dotsb,C_{r,(0)}+D_{r,1}\hat{K}_{1}(P\\ +Z)+D_{r,2}L+\hat{\varTheta}\hat{E}_{r})$ associated with the system $(A_{L}+B_{1}K_{L}(\tilde{P}_{L}),C_{1L}+D_{1,1}K_{L}(\tilde{P}_{L}),\dotsb,C_{rL}+D_{r,1}K_{L}(\tilde{P}_{L}))$ is mean-square stable, which implies the system $\hat{E}_{0},\hat{E}_{1},\dotsb,\hat{E}_{r};A_{(0)}+B_{1}\hat{K}_{1}(P+Z)+B_{2}L,C_{1,(0)}+D_{1,1}\hat{K}_{1}(P+Z)+D_{1,2}L,\dotsb,C_{r,(0)}+D_{r,1}\hat{K}_{1}(P+Z)+D_{r,2}L]$ is stochastically detectable.
@@
By Lemma 3.5, we have $\text{Spec}\,\mathcal{L}^{*}_{P+Z}\subset\mathbb{C}^{-}$. Thus, the proof of Conclusion (ii.) is completed. ∎

### v2

@@ 3.2 Existence and Uniqueness of Stabilizing Solution for ARE with Sign-Definite Quadratic Term
  • $0\in\bm{\Gamma}^{[A,C_{1},\dotsb,C_{r};B,D_{1},\dotsb,D_{r};Q,R]}$.
  • The system $A,C_{1},\dotsb,C_{r};\ B,D_{1},\dotsb,D_{r}]$ is mean-square stabilizable.
  • There exists a matrix set $\{E_{0},E_{1},\dotsb,E_{r}\}$ satisfying $\sum_{l=0}^{r}E_{l}^{\top}E_{l}=Q$ such that the system
    $E_{0},E_{1},\dotsb,E_{r};A,C_{1},\dotsb,C_{r}]$
@@
By Theorem 4.7 in [Drgan2004], along with the condition $0\in\bm{\Gamma}$ and the mean-square stabilizability of $\left[A,C_{1},\dotsb,C_{r};B,D_{1},\dotsb,D_{r}\right]$, the ARE (13) has a maximal solution $P_{\text{max}}\in\overline{\mathbb{S}}_{+}^{n}$ such that $R+\sum_{l=1}^{r}D_{l}^{\top}P_{\text{max}}D_{l}\succ 0$.
Using Proposition 3.2, the ARE for $P_{\text{max}}$ can be rewritten as
@@
Since the system $\left[E_{0},E_{1},\dotsb,E_{r};A,C_{1},\dotsb,C_{r}\right]$ is stochastically detectable, there exists $\varTheta\in\mathbb{R}^{n\times q}$ such that $\left(A+\varTheta E_{0},C_{1}+\varTheta E_{1},\dotsb,C_{r}+\varTheta E_{r}\right)$ is mean-square stable. Let $\hat{\varTheta}\in\mathbb{R}^{n\times[(q+m)(r+1)]}$ and $\hat{\varTheta}$=
$\begin{bmatrix}\varTheta^{\times(r+1)}&-B\sqrt{(r+1)R^{-1}}&-D_{1}\sqrt{(r+1)R^{-1}}&\dotsb&-D_{r}\sqrt{(r+1)R^{-1}}\end{bmatrix}$,
then the system
$\left(A+BT(P_{\text{max}})+\hat{\varTheta}\hat{E}_{0},C_{1}+D_{1}T(P_{\text{max}})+\hat{\varTheta}\hat{E}_{1},\dotsb,C_{r}+D_{r}T(P_{\text{max}})+\hat{\varTheta}\hat{E}_{r}\right)$
is mean-square stable.
By Lemma 3.5, we get the system
$(A+BT(P_{\text{max}}),\ C_{1}+D_{1}T(P_{\text{max}}),\dotsb,C_{r}+D_{r}T(P_{\text{max}}))$
is mean-square stable. Hence, $P_{\text{max}}$ is a stabilizing solution of ARE (13). By Theorem 5.6.5 in [Dragan2013book], the stabilizing solution is unique, so ARE (13) has a unique stabilizing solution $P\in\overline{\mathbb{S}}_{+}^{n}$ such that $R+\sum_{l=1}^{r}D_{l}^{\top}PD_{l}\succ 0$. ∎
@@ 3.3 Convergence Analysis of Iterative Sequence
As a foundation for the convergence analysis, we define two linear operators associated with the interrelated iterative steps. Lemma 3.8 clarifies their essential properties, which are crucial because they collectively ensure the boundedness of the iteration sequence. This guarantee of boundedness is then directly utilized to prove convergence in Theorem 3.9.
@@
$\mathcal{L}^{*}_{P}(Y)=Y(A_{(0)}+B_{1}\hat{K}_{1}(P)+B_{2}L)+(A_{(0)}+B_{1}\hat{K}_{1}(P)+B_{2}L)^{\top}Y+$ (14)
$\sum_{l=1}^{r}(C_{l,(0)}+D_{l,1}\hat{K}_{1}(P)+D_{l,2}L)^{\top}Y(C_{l,(0)}+D_{l,1}\hat{K}_{1}(P)+D_{l,2}L),\,\forall\,Y\in\mathbb{S}^{n};$
$\mathcal{L}^{*}_{P+Z}(Y)=Y(A_{(0)}+B_{1}\hat{K}_{1}(P+Z)+B_{2}L)+(A_{(0)}+B_{1}\hat{K}_{1}(P+Z)+B_{2}L)^{\top}Y+$ (15)
$\sum_{l=1}^{r}(C_{l,(0)}+D_{l,1}\hat{K}_{1}(P+Z)+D_{l,2}L)^{\top}Y(C_{l,(0)}+D_{l,1}\hat{K}_{1}(P+Z)+D_{l,2}L),\,\forall\,Y\in\mathbb{S}^{n},$
where $A_{(0)},C_{l,(0)}(1\leq l\leq r)$ are defined in (9) and
@@
Assume $0\in\mathrm{Dom}\,\mathcal{G}$ and that there exists an $L\in\mathcal{A}$. Let $P,Z\in\mathbb{S}^{n}$ with the following properties:
@@
    $\quad\times[B_{2}^{\top}Z+\sum_{l=1}^{r}D_{l,2}^{\top}Z(C_{l,(0)}+D_{l,1}\hat{K}_{1}(P)+D_{l,2}\hat{K}_{2}(P))].$
If $\mathcal{L}^{*}_{P}$, $\mathcal{L}^{*}_{P+Z}$ are linear operators associated with the triple $(P,Z,L)$ via (14) and (15) and $\tilde{P}_{L}$ is the stabilizing solution of the ARE (5) associated with $L$, then the following assertions hold:
(i). If   $\operatorname{Spec}\,\mathcal{L}^{*}_{P}\subset\mathbb{C}^{-}$, then $\tilde{P}_{L}\succeq P+Z$.
(ii). If   $\tilde{P}_{L}\succeq P+Z$, then $\operatorname{Spec}\,\mathcal{L}^{*}_{P+Z}\subset\mathbb{C}^{-}$.
@@
First, we prove the (i). From Proposition 3.2 and Proposition 3.3, we can obtain:
@@
where $J(P)=\begin{bmatrix}\hat{K}_{1}(P)\\ L\end{bmatrix}$, $\mathcal{L}^{*}_{P}$ is defined in (14), $\hat{K}_{1}(P)$ and $\hat{K}(P)$ is defined in (16).
Combining (17) and Proposition 3.3, we find that $P+Z$ solves the following ARE:
@@
where $\hat{N}_{1}(P,Z)=B_{1}^{\top}Z+\sum_{l=1}^{r}D_{l,1}^{\top}Z(C_{l,(0)}+D_{l,1}\hat{K}_{1}(P)+D_{l,2}\hat{K}_{2}(P))$, $\hat{N}_{2}(P,Z)=B_{2}^{\top}Z+\sum_{l=1}^{r}D_{l,2}^{\top}Z\\ (C_{l,(0)}+D_{l,1}\hat{K}_{1}(P)+D_{l,2}\hat{K}_{2}(P))$ and $\mathbb{R}_{22}^{\sharp}(P+Z)$ is defined in (10). Moreover,
$(J(P)-\hat{K}(P+Z))^{\top}R(P+Z)(J(P)-\hat{K}(P+Z))=$ (19)
$(\hat{K}_{1}(P)-\hat{K}_{1}(P+Z))^{\top}\mathbb{R}_{22}^{\sharp}(P+Z)(\hat{K}_{1}(P)-\hat{K}_{1}(P+Z))+H^{\top}_{1}H_{1}=$
$H^{\top}_{1}H_{1}+(\hat{N}_{1}(P,Z)-R_{12}(P+Z)R_{22}(P+Z)^{-1}\hat{N}_{2}(P,Z))^{\top}\mathbb{R}_{22}^{\sharp}(P+Z)^{-1}$
$\quad\times(\hat{N}_{1}(P,Z)-R_{12}(P+Z)R_{22}(P+Z)^{-1}\hat{N}_{2}(P,Z)),$
where $H_{1}=\begin{bmatrix}R_{22}(P+Z)^{-\frac{1}{2}}R_{21}(P+Z)&R_{22}(P+Z)^{\frac{1}{2}}\end{bmatrix}(J(P)-\hat{K}(P+Z))$.
Next, by virtue of Proposition 3.2, $\tilde{P}_{L}$ can be rewritten as:
@@
$-(\hat{K}_{1}(P)-K_{L}(\tilde{P}_{L}))^{\top}(R_{11}+\sum_{l=1}^{r}D_{l,1}^{\top}\tilde{P}_{L}D_{l,1})(\hat{K}_{1}(P)-K_{L}(\tilde{P}_{L}))=0,$
where $K_{L}(\tilde{P}_{L})=-(R_{11}+\sum_{l=1}^{r}D_{l,1}^{\top}\tilde{P}_{L}D_{l,1})^{-1}(B^{\top}_{1}\tilde{P}_{L}+\sum_{l=1}^{r}D^{\top}_{l,1}\tilde{P}_{L}C_{l}+R_{12}L)$.
Subtracting (18) from (20) and combining it with (19) yields:
@@
Since the $\operatorname{Spec}\,\mathcal{L}^{*}_{P}\subset\mathbb{C}^{-}$ and $H^{\top}H\succeq 0$, by a similarly proof of (a) $\Rightarrow$ (b) in Lemma 3.5, we have $\tilde{P}_{L}\succeq P+Z$. This completes the proof of (i).
We now prove (ii). Similarly, applying Proposition 3.2, Proposition 3.3 and combining (17), we find that $P+Z$ solves the following ARE:
$\mathcal{L}^{*}_{P+Z}(P+Z)+\hat{K}^{\top}_{1}(P+Z)R_{12}L+L^{\top}R_{21}\hat{K}_{1}(P+Z)+\hat{K}_{1}^{\top}(P+Z)R_{11}\hat{K}_{1}(P+Z)$ (21)
$+Q_{L}-(J(P+Z)-\hat{K}(P+Z))^{\top}R(P+Z)(J(P+Z)-\hat{K}(P+Z))$
$+(\hat{N}_{1}(P,Z)-R_{12}(P+Z)R_{22}(P+Z)^{-1}\hat{N}_{2}(P,Z))^{\top}\mathbb{R}_{22}^{\sharp}(P+Z)^{-1}$
@@
$\mathcal{L}^{*}_{P+Z}(\tilde{P}_{L})+\hat{K}^{\top}_{1}(P+Z)R_{12}L+L^{\top}R_{21}\hat{K}_{1}(P+Z)+\hat{K}_{1}^{\top}(P+Z)R_{11}\hat{K}_{1}(P+Z)+$ (22)
$Q_{L}-(\hat{K}_{1}(P+Z)-K_{L}(\tilde{P}_{L}))^{\top}(R_{11}+\sum_{l=1}^{r}D_{l,1}^{\top}\tilde{P}_{L}D_{l,1})(\hat{K}_{1}(P+Z)-K_{L}(\tilde{P}_{L}))=0,$
@@
$\begin{aligned} (\hat{K}_{1}(P+Z)-\hat{K}_{1}(P+Z))^{\top}\mathbb{R}_{22}^{\sharp}(P+Z)(\hat{K}_{1}(P+Z)-\hat{K}_{1}(P+Z))+H^{\top}_{2}H_{2}\end{aligned},$ (23)
where $H_{2}=\begin{bmatrix}R_{22}(P+Z)^{-\frac{1}{2}}R_{21}(P+Z)&R_{22}(P+Z)^{\frac{1}{2}}\end{bmatrix}(J(P+Z)-\hat{K}(P+Z))$.
Subtracting (21) from (22) and combining it with (23) yields:
$\mathcal{L}^{*}_{P+Z}(\tilde{P}_{L}-P-Z)+H_{2}^{\top}H_{2}-(\hat{K}_{1}(P+Z)-K_{L}(\tilde{P}_{L}))^{\top}(R_{11}+\sum_{l=1}^{r}D_{l,1}^{\top}\tilde{P}_{L}D_{l,1})$
$\times(\hat{K}_{1}(P+Z)-K_{L}(\tilde{P}_{L}))-(\hat{N}_{1}(P,Z)-R_{12}(P+Z)R_{22}(P+Z)^{-1}\hat{N}_{2}(P,Z))^{\top}$
$\times\mathbb{R}_{22}^{\sharp}(P+Z)^{-1}(\hat{N}_{1}(P,Z)-R_{12}(P+Z)R_{22}(P+Z)^{-1}\hat{N}_{2}(P,Z))=0.$
@@
Let $\Delta=\tilde{P}_{L}-P-Z$, we have: $\mathcal{L}^{*}_{P+Z}\Delta+\sum_{l=0}^{r}\hat{E}_{l}^{\top}\hat{E}_{l}=0,\Delta\succeq 0$, where $\hat{E}_{l}\in\mathbb{R}^{[m_{2}+m_{1}(r+2)]\times n}(0\leq l\leq r)$,
@@
$\hat{E}_{l}=\begin{bmatrix}\mathbb{O}_{m_{1}\times n}^{\times(l)}\\ \sqrt{-\frac{1}{r+1}(R_{11}+\sum_{l=1}^{r}D_{l,1}^{\top}\tilde{P}_{L}D_{l,1})}(\hat{K}_{1}(P+Z)-K_{L}(\tilde{P}_{L}))\\ \mathbb{O}_{m_{1}\times n}^{\times(r-l)}\\ \frac{1}{\sqrt{r+1}}H_{2}\\ \sqrt{-\frac{1}{r+1}\mathbb{R}_{22}^{\sharp}(P+Z)^{-1}}(\hat{N}_{1}(P,Z)-R_{12}(P+Z)R_{22}(P+Z)^{-1}\hat{N}_{2}(P,Z))\end{bmatrix}$.
Choose $\hat{\varTheta}\in\mathbb{R}^{n\times[m_{2}+m_{1}(r+2)]}$, and
@@
$\hat{\varTheta}=\begin{bmatrix}-\sqrt{-(r+1)(R_{11}+\sum_{l=1}^{r}D_{l,1}^{\top}\tilde{P}_{L}D_{l,1})^{-1}}^{\top}B_{1}^{\top}\\ -\sqrt{-(r+1)(R_{11}+\sum_{l=1}^{r}D_{l,1}^{\top}\tilde{P}_{L}D_{l,1})^{-1}}^{\top}D_{1,1}^{\top}\\ \dotsb\\ -\sqrt{-(r+1)(R_{11}+\sum_{l=1}^{r}D_{l,1}^{\top}\tilde{P}_{L}D_{l})^{-1}}^{\top}D_{r,1}^{\top}\\ \mathbb{O}_{n\times m_{2}}^{\top}\\ \mathbb{O}_{n\times m_{1}}^{\top}\end{bmatrix}^{\top}$.
@@
We obtain that $(A_{(0)}+B_{1}\hat{K}_{1}(P+Z)+B_{2}L+\hat{\varTheta}\hat{E}_{0},C_{1,(0)}+D_{1,1}\hat{K}_{1}(P+Z)+D_{1,2}L+\hat{\varTheta}\hat{E}_{1},\dotsb,C_{r,(0)}+D_{r,1}\hat{K}_{1}(P+Z)+D_{r,2}L+\hat{\varTheta}\hat{E}_{r})$ associated with the system $(A_{L}+B_{1}K_{L}(\tilde{P}_{L}),C_{1L}+D_{1,1}K_{L}(\tilde{P}_{L}),\dotsb,C_{rL}+D_{r,1}K_{L}(\tilde{P}_{L}))$ is mean-square stable.
@@
This implies the system $\hat{E}_{0},\hat{E}_{1},\dotsb,\hat{E}_{r};A_{(0)}+B_{1}\hat{K}_{1}(P+Z)+B_{2}L,C_{1,(0)}+D_{1,1}\hat{K}_{1}(P+Z)+D_{1,2}L,\dotsb,\\ C_{r,(0)}+D_{r,1}\hat{K}_{1}(P+Z)+D_{r,2}L]$ is stochastically detectable. By Lemma 3.5, we have $\operatorname{Spec}\,\mathcal{L}^{*}_{P+Z}\subset\mathbb{C}^{-}$. Thus, the proof of (ii). is completed. ∎